# Parameter-Efficient Fine-Tuning for Large Models: A Comprehensive Survey

### v5

@@ III-A2 Soft Prompt
Alternatively, prompt tuning presents an additional approach for refining the model to achieve improved performance through fine-tuning. Instead of optimizing discrete token representations through in-context learning, there is a prevailing belief that the continuous embedding space of soft prompts inherently contains more information [96]. Drawing inspiration from this concept, researchers directly append adjustable vectors, referred to as soft prompts, to the start of the input sequence. This can be represented as follows:
@@
Prefix-tuning [35] introduces learnable vectors that are prepended to keys $k$ and values $v$ across all Transformer layers. To ensure stability during the optimization process, Prefix-tuning adopts a reparameterization strategy, which utilizes an MLP layer to generate these prefix vectors rather than optimizing them directly. After fine-tuning, only the prefix vectors are saved for inference. This technique is adapted and improved in several studies [36, 37, 38]. For instance, p-tuning v2 [37] removes reparameterization and expands its usage to broader model scales and NLP tasks. APT (Adaptive Prefix Tuning) [38] enhances Prefix-tuning by introducing adaptive gate mechanism to control the prefix importance in each layer. Concurrent work p-tuning [39] and prompt-tuning [40] apply learnable vectors only at the initial word embedding layer rather than all layers to enhance training and inference efficiency. It’s important to highlight that prompt-tuning demonstrates its effectiveness primarily in the context of large models, specifically those with over 11 billion parameters [40]. Complementing this, Xprompt [41] eliminates the negative prompt tokens through a hierarchical structured pruning, which closes the performance gap at smaller model scales. The work by [97] provides some theoretical analysis towards prompt tuning, demonstrating its universality and limitations in limited-depth Transformers. IDPG (Instance-Dependent Prompt Generation) [42] improves prompt tuning by generating prompts based on each input sentence with a lightweight prompt generator. In a related approach, LPT (Late Prompt Tuning) [43] also leverages a prompt generator to obtain instance-aware prompt. Unlike previous work, LPT adds these prompts only after an intermediate layer, rather than at the initial or all layers. This strategic placement eliminates the gradient calculation below the intermediate layer, thereby significantly accelerating the training speed. Simultaneously, LPT can improve the overall performance due to the shorter backpropagation path preserves more task-related information. Inspired by LPT, SPT (Selective Prompt Tuning) [44] delves deeper into the importance of prompt inserting strategies. It introduces a learnable probabilistic gate in each layer to determine whether to use the prompt propagated from the previous layer or inject a newly generated prompt. APrompt [45] employs another prompt inserting strategy. In addition to input prompts inserted at the beginning of the input sequence for each Transformer layer, APrompt also prepends additional learnable prompts to the respective query, key, and value matrices in the self-attention blocks to learn new attention patterns. Besides, APrompt incorporates the learning of a task-specific head.
@@
The concept of soft prompts has been employed for various downstream tasks [98, 99], although their training can be prone to instability and slow convergence. To address this, SPoT [46] uses a source prompt learned from one or multiple tasks to initialize prompts for new tasks. Similarly, transfer of soft prompts from one task to initialize another is proposed in TPT (transferable prompt tuning) [47], which demonstrates that a better prompt initialization results in a large training convergence speedup. InfoPrompt [48] develops two mutual information based loss functions, i.e., head loss and representation loss, to find better prompt initialization and learn sufficient task-relevant information, thereby also expediting convergence. PTP [49] delves into the root causes of training instability. It identifies the steep nature of the loss landscape in conventional prompt tuning, where minor variations in input data can lead to significant loss fluctuations. To mitigate this, PTP introduces perturbation-based regularizers to smooth the loss landscape and consequently stabilize the training process. DePT [52] decomposes the soft prompt into a shorter soft prompt with a pair of low-rank matrices, which are optimised with two distinct learning rates. This strategy not only improves the performance but also enhances training and inference efficiency. SMoP (Sparse Mixture-of-Prompts) [51] reduce the training and inference cost by utilizing short soft prompts. During training, multiple short soft prompts are trained, each tailored to specific subsets of the dataset. During inference, SMoP integrates a gating mechanism that routes each input instance to an appropriate short prompt. This technique not only increases efficiency in both training and inference stages but also retains performance comparable to those achieved with longer soft prompts. To further cut down the number of soft prompt parameters, IPT (Intrinsic Prompt Tuning) [50] identifies an intrinsic task subspace by training an auto-encoder on multiple tasks. Tuning on new tasks then requires adjusting only a few parameters within this subspace, significantly reducing the number of training parameters.
@@ III-B Selective PEFT
Diff pruning [57] is a representative work that applies a learnable binary mask to the model weights during fine-tuning. To achieve parameter efficiency, the mask is regularized by a differentiable approximation of the $L_{0}$-norm penalty. PaFi [59] simply select model parameters with the smallest absolute magnitude as trainable. FishMask [60] determines parameter importance using the approximate Fisher information. It then selects the top k parameters based on this information to form the mask $M$. Similarly, Fish-Dip [61] also uses Fisher information to calculate $M$, but the mask will be re-calculated dynamically in each train period. LT-SFT [62] introduces another technique to determines parameter importance inspired by the Lottery Ticket Hypothesis [100, 101], where the the subset of parameters that change the most during an initial fine-tuning stage is selected to form the mask $M$. SAM [63] proposes a second-order approximation method, which approximates the original problem with an analytically solvable optimization function, to help decide the parameter mask. Child-tuning [64] proposes two approaches to select a child network during each training iteration, where only the parameters within this child network can be updated.
Figure 7: Illustration of how two parameter masking methods. Blue represents frozen, while yellow represents trainable.
@@
However, above unstructured parameter masking results in an uneven distribution of non-zero masks and diminished hardware efficiency when implementing PEFT. As shown in Figure 7, the structured mask organize parameter masking in regular patterns, unlike unstructured ones that apply it randomly, thus can enhances computational and hardware efficiency during training. Therefore, various structured selective PEFT techniques have undergone extensive investigation. Diff pruning proposes a structured pruning strategy by partitioning the weight parameters into local groups and strategically eliminating them together. Similarly, FAR [65] fine-tunes BERT models by grouping weights of the FFN in Transformer blocks into nodes, then rank and select the learner nodes using $L_{1}$ norm. To further reduce the memory access frequency, they also reconfigure the FFN by grouping the learner nodes together. Bitfit [66] is proposed to only fine-tunes the bias parameters of each DNN layer, and achieves competitive results for small models. However, this method fails to handle large models. The work by [58] applies NAS to Bitfit, where S-BitFit keeps the structural nature in Bitfit that restrict NAS algorithm must choose whether $\delta b=0$ or not for each bias module. Similar to Bitfit that fine-tunes a specific module in Transformer, Xattn Tuning [67] fine-tunes only the cross-attention layers. SPT (sensitivity-aware visual parameter-efficient fine-tuning) [68] first identifies the sensitive parameters measured by the loss reduction when being tuned. This sensitivity is calculated using a first-order Taylor expansion, derived from a single forward and backward pass before fine-tuning in oneshot. Next, SPT finds the weight matrices whose number of sensitive parameters exceeds a pre-defined threshold, and then applies a selected PEFT technique (e.g., LoRA and Adapter) to these targeted weights to achieve structural tuning.
@@ III-C Reparameterized PEFT
Earlier research studies [69] have shown that common pre-trained models exhibit an exceptionally low intrinsic dimensionality. In other words, it is possible to find a low-dimensional reparameterization that is effective for fine-tuning as the entire parameter space. Intrinsic SAID [69] is the pioneering work in investigating the intrinsic dimension feature during the fine-tuning of LLMs. However, the most widely recognized reparameterization technique is LoRA (Low Rank Adaptation) [70, 102], as shown in Figure 8 (a). For a given pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA introduces two trainable weight matrices, $W_{\text{up}}\in\mathbb{R}^{d\times r}$ and $W_{\text{down}}\in\mathbb{R}^{r\times k}$ where the rank $r\ll min(d,k)$, operating in parallel to $W_{0}$. Let $h_{in}$ represent the input. Under normal conditions, the output through $W_{0}$ is $h_{out}=W_{0}h_{in}$. Instead, LoRA modifies this output by introducing an incremental update $\Delta W$ that encapsulates task-specific knowledge:
@@
In LoRA training, selecting an appropriate rank has always been a challenging issue. To address this, DyLoRA [76], as depicted in Figure 8 (b), trains LoRA module on a range of ranks within a predefined training budget, rather than adhering to a single, fixed rank. Specifically, for a given rank range $R=\{r_{\text{min}},r_{\text{min}}+1,\ldots,r_{\text{max}}\}$, DyLoRA dynamically chooses a rank $r\in R$ at each iteration of the training process. Consequently, the matrices $W_{\text{down}}$ and $W_{\text{up}}$ are tailored for the selected rank $r$, resulting in truncated versions $W_{\text{down}\downarrow r}=W_{\text{down}}[1:r,:]$ and $W_{\text{up}\downarrow r}=W_{\text{up}}[:,1:r]$, and the subsequent forward and backward pass during this iteration will be restricted on $W_{\text{down}\downarrow r}$ and $W_{\text{up}\downarrow r}$ instead of $W_{\text{down}}$ and $W_{\text{up}}$. With this dynamic and search-free approach, DyLoRA significantly reduces the training time required to find an optimal and fixed LoRA rank for specific tasks. AdaLoRA [77] reformulates the $\Delta W$ with a singular value decomposition (SVD), denoted as $\Delta W=P\Lambda Q$, where $P\in\mathbb{R}^{d\times r}\text{ and }Q\in\mathbb{R}^{r\times k}$ are orthometric, $\Lambda$ is a diagonal matrix containing sigular values $\{\lambda_{i}\}_{1\leq i\leq r}$. All the three weight matrices are made learnable. During training, the singular values are pruned iteratively based on their importance scores, which are constructed from moving average of the magnitude of gradient-weight product. To ensure the orthogonality between $P\text{ and }Q$, i.e., $P^{T}P=QQ^{T}=I$, an additional regularizer term is included in the loss:
@@
This adaptive approach enables the model to dynamically adjust the rank within each LoRA module, effectively managing its parameter counts based on the significance of the weight matrices. However, according to SoRA [78], the importance scores used in AdaLoRA is heuristically constructed, which lacks rigorous theoretical motivation. Additionally, both moving average operation and calculation of Eq. 13 introduces extra computation cost during training. To address this, SoRA eliminates the orthogonality premise of $P$ and $Q$. Instead, a gating unit $g\in\mathbb{R}^{r}$ between $W_{\text{up}}$ and $W_{\text{down}}$ is directly applied and optimized:
@@
where $\odot$ is Hadamard product. The gate $g$ is updated using a variation of proximal gradient iteration for $l_{1}$ loss [103, 104], which has a clear mathematical meaning and do not need the heuristic premise. After training, the zeroed-out gate units are pruned by removing the corresponding columns and rows in $W_{\text{down}}$ and $W_{\text{up}}$.
Several subsequent studies have aimed to improve LoRA’s performance in various aspects. For instance, Laplace-LoRA [81] notices that fine-tuned LLMs often exhibit overconfidence. To enhance the calibration of fine-tuned LLMs, Laplace-LoRA utilizes a Bayesian approach, specifically a post-hoc Laplace approximation [105, 106], to the posterior over the LoRA parameters. LoRA Dropout [82] introduces random noises to the learnable low-rank matrices and increases parameter sparsity to reduce the risk of overfitting. LoRA+ [84] proposes to set different learning rates for the LoRA matrices $W_{\text{down}}$ and $W_{\text{up}}$, such that $\eta_{\text{up}}=\lambda\eta_{\text{down}}$ with $\lambda>1$ fixed and tune $\eta_{\text{down}}$.
Thanks to the modular design of LoRA, many studies incorporate multiple LoRA modules in their frameworks to enhance performance. For example, LoRAHub aggregates various LoRA modules trained on different tasks. Given a handful of examples from a new task, LoRAHub can autonomously compose compatible LoRA modules without human intervention via a gradient-free method Shiwa [107]. MOELoRA employs a Mixture-of-Experts (MOE) approach to train LoRA in a multi-task setting, resulting in multiple expert LoRA modules. To retrieve parameters for certain task, MOELoRA utilizes a task-motivated gate function that assigns contribution weights to each expert based on the task ID, and the final parameters is calculated through a weighted sum of all experts.
@@ III-D Hybrid PEFT
The efficacy of various PEFT methods can significantly differ across different tasks. As a result, numerous studies aim to either combine the advantages of diverse PEFT approaches or seek to establish a unified perspective by analyzing the similarities among these methods. For instance, UniPELT[90] integrates LoRA, prefix-tuning, and adapters into each Transformer block. To control which PEFT submodules should be activated, they also introduce a gating mechanism. This mechanism consists of three small FFNs that each produce a scalar value $\mathcal{G}\in(0,1)$, which is then applied to the LoRA, prefix, and adapter matrices, respectively. Across various setups, UniPELT has consistently shown improvements in accuracy ranging from 1% to 4%. S4 [91] explores design spaces for several PEFT methods (i.e., Adapter (A), Prefix (P), BitFit (B), and LoRA (L)) to uncover underlying design patterns. After a series experiments, their findings include: (1) Applying the spindle grouping partitioning for Transformer layers, which results in four layer groups $G_{i}$ for $i\in\{1\dots 4\}$. Layers in one group have similar behaviors together, which means should be apply similar PEFT strategies. (2) Allocating the number of trainable parameters to layers uniformly. (3) Tuning all the groups. (4) Assigning different PEFT strategies in different group. The resulting design space that has the best performance is:
@@
MAM Adapter[26] explores the intrinsic similarity between three additive PEFT methods: adapters, prefix-tuning, and LoRA, which leads to the development of three variants: Parallel Adapter, which places adapter layers alongside specific layers (SA or FFN) instead of after them; Multi-head Parallel Adapter, which divides the parallel adapter into multiple heads, each affecting the head attention output in SA; and Scaled Parallel Adapter, which adds a scaling term after the parallel adapter layer, similar to LoRA. Extensive experimentation revealed that the most effective configuration involves using prefix-tuning in the SA layer and the scaled parallel adapter in the FFN layer, which is called MAM Adapter. LLM-Adapters [94] builds an easy-to-use framework that incorporates various PEFT techniques into LLMs. Through comprehensive benchmarking across multiple datasets, the study reveals several key insights: (1) The most effective locations for series adapters, parallel adapters, and LoRA are after the MLP layers, alongside the MLP layers, and simultaneously following the Attention layers and MLP layers, respectively. (2) Smaller LLMs utilizing PEFT can achieve competitive or even superior results on certain tasks when compared to their larger counterparts. (3) With appropriate in-distribution fine-tuning data, smaller models are capable of surpassing larger models in task-specific performance.
Several studies leverage neural architecture search (NAS) to find better PEFT combination approaches. For example, NOAH [92] discovers that different PEFT configurations are specifically tailored for different tasks. To address this issue, NOAH employs NAS to identify the most effective PEFT configurations for each dataset. Specifically, NOAH’s searching space encompasses three PEFT methods: Adapter, LoRA, and Visual Prompt Tuning (VPT). It utilizes AutoFormer [108], a one-shot NAS algorithm, for the efficient discovery of optimal prompt modules. In a related vein, AUTOPEFT [93] first establishes a searching space that includes serial adapters, parallel adapters, and prefix tuning. After that, they proposes an effective NAS methods based on a high-dimensional multi-dimensional Bayesian optimisation [109]. Both NOAH and AUTOPEFT demonstrate the capability of NAS in enhancing PEFT configurations across a variety of tasks.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, node options=align=center, rounded corners, align=center, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.0em,font=, where level=2text width=5.6em,font=, where level=3text width=6.8em,font=, [ Efficient PEFT Design, ver [ PEFT Pruning [ AdapterDrop [110], SparseAdapter [111], SPLoRA [112], LoRAPruning [113], ProPETL [114] , leaf, text width=38.4em ] ] [ PEFT

### v7

@@ III-A2 Soft Prompt
Alternatively, prompt tuning presents an additional approach for refining the model to achieve improved performance through fine-tuning. Instead of optimizing discrete token representations through in-context learning, there is a prevailing belief that the continuous embedding space of soft prompts inherently contains more information [103]. Drawing inspiration from this concept, researchers directly prepend adjustable vectors, referred to as soft prompts, to the start of the input sequence. This can be represented as follows:
@@
Prefix-tuning [41] introduces learnable vectors that are prepended to keys $k$ and values $v$ across all Transformer layers. To ensure stability during the optimization process, Prefix-tuning adopts a reparameterization strategy, which utilizes an MLP layer to generate these prefix vectors rather than optimizing them directly. After fine-tuning, only the prefix vectors are saved for inference. This technique has been adapted and improved in several studies [42, 43, 44]. For instance, p-tuning v2 [43] removes reparameterization and expands its usage to broader model scales and NLP tasks. APT (Adaptive Prefix Tuning) [44] enhances Prefix-tuning by introducing an adaptive gate mechanism to control the prefix importance in each layer. Concurrent work p-tuning [45] and prompt-tuning [46] apply learnable vectors only at the initial word embedding layer rather than all layers to enhance training and inference efficiency. It’s important to highlight that prompt-tuning demonstrates its effectiveness primarily in the context of large models, specifically those with over 11 billion parameters [46]. Complementing this, Xprompt [47] eliminates the negative prompt tokens through a hierarchically structured pruning, which closes the performance gap at smaller model scales. [104] provides some theoretical analysis towards prompt tuning, demonstrating its universality and limitations in limited-depth Transformers. IDPG (Instance-Dependent Prompt Generation) [48] improves prompt tuning by generating prompts based on each input sentence with a lightweight prompt generator. In a related approach, LPT (Late Prompt Tuning) [49] also leverages a prompt generator to obtain instance-aware prompt. Unlike previous work, LPT adds these prompts only after an intermediate layer, rather than at the initial or all layers. This strategic placement eliminates the gradient calculation below the intermediate layer, thereby significantly accelerating the training speed. Simultaneously, LPT can improve the overall performance due to the shorter backpropagation path preserves more task-related information. Inspired by LPT, SPT (Selective Prompt Tuning) [50] delves deeper into the importance of prompt inserting strategies. It introduces a learnable probabilistic gate in each layer to determine whether to use the prompt propagated from the previous layer or inject a newly generated prompt. APrompt [51] employs another prompt inserting strategy. In addition to input prompts inserted at the beginning of the input sequence for each Transformer layer, APrompt also prepends additional learnable prompts to the respective query, key, and value matrices in the self-attention blocks to learn new attention patterns. Besides, APrompt incorporates the learning of a task-specific head.
@@
The concept of soft prompts has been employed for various downstream tasks [105, 106], although their training can be prone to instability and slow convergence. To address this, SPoT [52] uses a source prompt learned from one or multiple tasks to initialize prompts for new tasks. Similarly, the transfer of soft prompts from one task to initialize another is proposed in TPT (transferable prompt tuning) [53], which demonstrates that a better prompt initialization results in a large training convergence speedup. InfoPrompt [54] develops two mutual information-based loss functions, i.e., head loss and representation loss, to find better prompt initialization and learn sufficient task-relevant information, thereby also expediting convergence. PTP [55] delves into the root causes of training instability. It identifies the steep nature of the loss landscape in conventional prompt tuning, where minor variations in input data can lead to significant loss fluctuations. To mitigate this, PTP introduces perturbation-based regularizers to smooth the loss landscape and consequently stabilize the training process. DePT [58] decomposes the soft prompt into a shorter soft prompt with a pair of low-rank matrices, which are optimized with two distinct learning rates. This strategy not only improves performance but also enhances training and inference efficiency. SMoP (Sparse Mixture-of-Prompts) [57] reduce the training and inference cost by utilizing short soft prompts. During training, multiple short soft prompts are trained, each tailored to specific subsets of the dataset. During inference, SMoP integrates a gating mechanism that routes each input instance to an appropriate short prompt. This technique not only increases efficiency in both training and inference stages but also retains performance comparable to those achieved with longer soft prompts. To further cut down the number of soft prompt parameters, IPT (Intrinsic Prompt Tuning) [56] identifies an intrinsic task subspace by training an auto-encoder on multiple tasks. Tuning on new tasks then requires adjusting only a few parameters within this subspace, significantly reducing the number of training parameters.
@@ III-B Selective PEFT
Diff pruning [63] is a representative work that applies a learnable binary mask to the model weights during fine-tuning. To achieve parameter efficiency, the mask is regularized by a differentiable approximation of the $L_{0}$-norm penalty. PaFi [65] simply select model parameters with the smallest absolute magnitude as trainable. FishMask [66] determines parameter importance using the approximate Fisher information. It then selects the top k parameters based on this information to form the mask $M$. Similarly, Fish-Dip [67] also uses Fisher information to calculate $M$, but the mask will be re-calculated dynamically in each train period. LT-SFT [68] introduces another technique to determine parameter importance inspired by the Lottery Ticket Hypothesis [107, 108], where the subset of parameters that change the most during an initial fine-tuning stage is selected to form the mask $M$. SAM [69] proposes a second-order approximation method, which approximates the original problem with an analytically solvable optimization function, to help decide the parameter mask. Child-tuning [70] proposes two approaches to select a child network during each training iteration, where only the parameters within this child network can be updated.
Figure 7: Illustration of two parameter masking methods.
@@
However, the above unstructured parameter masking results in an uneven distribution of non-zero masks and diminished hardware efficiency when implementing PEFT. As shown in Figure 7, the structured mask organizes parameter masking in regular patterns, unlike unstructured ones that apply it randomly, thus enhancing computational and hardware efficiency during training. Therefore, various structured selective PEFT techniques have undergone extensive investigation. Diff pruning proposes a structured pruning strategy by partitioning the weight parameters into local groups and strategically eliminating them together. Similarly, FAR [71] fine-tunes BERT models by grouping weights of the FFN in Transformer blocks into nodes, then ranking and selecting the learner nodes using $L_{1}$ norm. To further reduce the memory access frequency, they also reconfigure the FFN by grouping the learner nodes. Bitfit [72] is proposed to only fine-tune the bias parameters of each DNN layer, and achieves competitive results for small models. However, this method fails to handle large models. [64] applies NAS to Bitfit, where S-BitFit keeps the structural nature in Bitfit that restricts NAS algorithm must choose whether $\delta b=0$ or not for each bias module. Similar to Bitfit that fine-tunes a specific module in Transformer, Xattn Tuning [73] fine-tunes only the cross-attention layers. SPT (sensitivity-aware visual parameter-efficient fine-tuning) [74] first identifies the sensitive parameters measured by the loss reduction when being tuned. This sensitivity is calculated using a first-order Taylor expansion, derived from a single forward and backward pass before fine-tuning in one shot. Next, SPT finds the weight matrices whose number of sensitive parameters exceeds a pre-defined threshold and then applies a selected PEFT technique (e.g., LoRA and Adapter) to these targeted weights to achieve structural tuning.
@@ III-C Reparameterized PEFT
Earlier research studies [75] have shown that common pre-trained models exhibit an exceptionally low intrinsic dimensionality. In other words, it is possible to find a low-dimensional reparameterization that is effective for fine-tuning as the entire parameter space. Intrinsic SAID [75] is the pioneering work in investigating the intrinsic dimension feature during the fine-tuning of LLMs. However, the most widely recognized reparameterization technique is LoRA (Low-Rank Adaptation) [76, 109], as shown in Figure 8 (a). For a given pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA introduces two trainable weight matrices, $W_{\text{up}}\in\mathbb{R}^{d\times r}$ and $W_{\text{down}}\in\mathbb{R}^{r\times k}$ where the rank $r\ll min(d,k)$, operating in parallel to $W_{0}$. Let $h_{in}$ represent the input. Under normal conditions, the output through $W_{0}$ is $h_{out}=W_{0}h_{in}$. Instead, LoRA modifies this output by introducing an incremental update $\Delta W$ that encapsulates task-specific knowledge:
@@
In LoRA training, selecting an appropriate rank has always been a challenging issue. To address this, DyLoRA [82], as depicted in Figure 8 (b), trains the LoRA module on a range of ranks within a predefined training budget, rather than adhering to a single, fixed rank. Specifically, for a given rank range $R=\{r_{\text{min}},r_{\text{min}}+1,\ldots,r_{\text{max}}\}$, DyLoRA dynamically chooses a rank $r\in R$ at each iteration of the training process. Consequently, the matrices $W_{\text{down}}$ and $W_{\text{up}}$ are tailored for the selected rank $r$, resulting in truncated versions $W_{\text{down}\downarrow r}=W_{\text{down}}[1:r,:]$ and $W_{\text{up}\downarrow r}=W_{\text{up}}[:,1:r]$, and the subsequent forward and backward pass during this iteration will be restricted on $W_{\text{down}\downarrow r}$ and $W_{\text{up}\downarrow r}$ instead of $W_{\text{down}}$ and $W_{\text{up}}$. With this dynamic and search-free approach, DyLoRA significantly reduces the training time required to find an optimal and fixed LoRA rank for specific tasks. AdaLoRA [83] reformulates the $\Delta W$ with a singular value decomposition (SVD), denoted as $\Delta W=P\Lambda Q$, where $P\in\mathbb{R}^{d\times r}\text{ and }Q\in\mathbb{R}^{r\times k}$ are orthometric, $\Lambda$ is a diagonal matrix containing singular values $\{\lambda_{i}\}_{1\leq i\leq r}$. All three weight matrices are made learnable. During training, the singular values are pruned iteratively based on their importance scores, which are constructed from the moving average of the magnitude of the gradient-weight product. To ensure the orthogonality between $P\text{ and }Q$, i.e., $P^{T}P=QQ^{T}=I$, an additional regularizer term is included in the loss:
@@
This adaptive approach enables the model to dynamically adjust the rank within each LoRA module, effectively managing its parameter counts based on the significance of the weight matrices. However, according to SoRA [84], the importance scores used in AdaLoRA are heuristically constructed, which lacks rigorous theoretical motivation. Additionally, both moving average operation and calculation of Eq. 13 introduce extra computation costs during training. To address this, SoRA eliminates the orthogonality premise of $P$ and $Q$. Instead, a gating unit $g\in\mathbb{R}^{r}$ between $W_{\text{up}}$ and $W_{\text{down}}$ is directly applied and optimized:
@@
where $\odot$ is Hadamard product. The gate $g$ is updated using a variation of proximal gradient iteration for $l_{1}$ loss [110, 111], which has a clear mathematical meaning and does not need the heuristic premise. After training, the zeroed-out gate units are pruned by removing the corresponding columns and rows in $W_{\text{down}}$ and $W_{\text{up}}$.
Several subsequent studies have aimed to improve LoRA’s performance in various aspects. For instance, Laplace-LoRA [87] notices that fine-tuned LLMs often exhibit overconfidence. To enhance the calibration of fine-tuned LLMs, Laplace-LoRA utilizes a Bayesian approach, specifically a post-hoc Laplace approximation [112, 113], to the posterior over the LoRA parameters. LoRA Dropout [88] introduces random noises to the learnable low-rank matrices and increases parameter sparsity to reduce the risk of overfitting. LoRA+ [90] proposes to set different learning rates for the LoRA matrices $W_{\text{down}}$ and $W_{\text{up}}$, such that $\eta_{\text{up}}=\lambda\eta_{\text{down}}$ with $\lambda>1$ fixed and tune $\eta_{\text{down}}$. MoSLoRA (Mixture-of-Subspaces LoRA) [91] decomposes LoRA into subspaces via structural reparameterization, then employs a learnable mixer, trained jointly with the original LoRA weights, to fuse the subspaces. Similarly to LoRA, MoSLoRA can also be merged into the original weights.
Thanks to the modular design of LoRA, many studies incorporate multiple LoRA modules in their frameworks to enhance performance. For example, LoRAHub aggregates various LoRA modules trained on different tasks. Given a handful of examples from a new task, LoRAHub can autonomously compose compatible LoRA modules without human intervention via a gradient-free method Shiwa [114]. MOELoRA employs a Mixture-of-Experts (MOE) approach to train LoRA in a multi-task setting, resulting in multiple expert LoRA modules. To retrieve parameters for certain tasks, MOELoRA utilizes a task-motivated gate function that assigns contribution weights to each expert based on the task ID, and the final parameters are calculated through a weighted sum of all experts.
@@ III-D Hybrid PEFT
The efficacy of various PEFT methods can significantly differ across different tasks. As a result, numerous studies aim to either combine the advantages of diverse PEFT approaches or seek to establish a unified perspective by analyzing the similarities among these methods. For instance, UniPELT [97] integrates LoRA, prefix-tuning, and adapters into each Transformer block. To control which PEFT submodules should be activated, they also introduce a gating mechanism. This mechanism consists of three small FFNs that each produce a scalar value $\mathcal{G}\in(0,1)$, which is then applied to the LoRA, prefix, and adapter matrices, respectively. Across various setups, UniPELT has consistently shown improvements in accuracy ranging from 1% to 4%. S4 [98] explores design spaces for several PEFT methods (i.e., Adapter (A), Prefix (P), BitFit (B), and LoRA (L)) to uncover underlying design patterns. After a series of experiments, their findings include: (1) Applying the spindle grouping partitioning for Transformer layers, which results in four layer groups $G_{i}$ for $i\in\{1\dots 4\}$. Layers in one group have similar behaviors together, which means should apply similar PEFT strategies. (2) Allocating the number of trainable parameters to layers uniformly. (3) Tuning all the groups. (4) Assigning different PEFT strategies in different groups. The resulting design space that has the best performance is:
@@
MAM Adapter[32] explores the intrinsic similarity between three additive PEFT methods: adapters, prefix-tuning, and LoRA, which leads to the development of three variants: Parallel Adapter, which places adapter layers alongside specific layers (SA or FFN) instead of after them; Multi-head Parallel Adapter, which divides the parallel adapter into multiple heads, each affecting the head attention output in SA; and Scaled Parallel Adapter, which adds a scaling term after the parallel adapter layer, similar to LoRA. Extensive experimentation revealed that the most effective configuration involves using prefix-tuning in the SA layer and the scaled parallel adapter in the FFN layer, which is called the MAM Adapter. LLM-Adapters [101] builds an easy-to-use framework that incorporates various PEFT techniques into LLMs. Through comprehensive benchmarking across multiple datasets, the study reveals several key insights: (1) The most effective locations for series adapters, parallel adapters, and LoRA are after the MLP layers, alongside the MLP layers, and simultaneously following the Attention layers and MLP layers, respectively. (2) Smaller LLMs utilizing PEFT can achieve competitive or even superior results on certain tasks when compared to their larger counterparts. (3) With appropriate in-distribution fine-tuning data, smaller models are capable of surpassing larger models in task-specific performance.
Several studies leverage neural architecture search (NAS) to find better PEFT combination approaches. For example, NOAH [99] discovers that different PEFT configurations are specifically tailored for different tasks. To address this issue, NOAH employs NAS to identify the most effective PEFT configurations for each dataset. Specifically, NOAH’s searching space encompasses three PEFT methods: Adapter, LoRA, and Visual Prompt Tuning (VPT). It utilizes AutoFormer [115], a one-shot NAS algorithm, for the efficient discovery of optimal prompt modules. In a related vein, AUTOPEFT [100] first establishes a searching space that includes serial adapters, parallel adapters, and prefix tuning. After that, they propose an effective NAS method based on a high-dimensional multi-dimensional Bayesian optimisation [116]. Both NOAH and AUTOPEFT demonstrate the capability of NAS in enhancing PEFT configurations across a variety of tasks.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, node options=align=center, rounded corners, align=center, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.0em,font=, where level=2text width=5.6em,font=, where level=3text width=6.8em,font=, [ Efficient PEFT Design, ver [ PEFT Pruning [ AdapterDrop [117], SparseAdapter [118], SPLoRA [119], LoRAPruning [120], ProPETL [121] , leaf, text width=38.4em, align = left ] ] [ PEFT